Consider the following tasks, times, and predecessors for an assembly of set top cable converter boxes:

TASK ELEMENT TIME
(MINUTES)
ELEMENT
PREDECESSOR
A 2
B 1 A
C 5 B
D 4 B
E 3 C, D
F 3 A
G 2 F
H 1 G
I 4 E, H

Given a cycle time of seven minutes, what is the efficiency of your layout? Use the longest task time as primary criterion with largest number of following tasks as secondary criterion. (Round your answer to the nearest whole percent.)
Answer #1

CYCLE TIME = 7

THEORETICAL NUMBER OF WORKSTATIONS REQUIRED = SIGMA(TASK TIMES) / CYCLE TIME

N = (2 + 1 + 5 + 4 + 3 + 3 + 2 + 1) / 7 = 4


BALANCED USING LONGEST TASK TIME AND GREATEST NUMBER OF FOLLOWING TIMES AS TIE BREAKER.

WORKSTATION

ASSIGN TASK

TASK TIME

TIME REMAINING

IDLE TIME

1

A

2

5

F

3

2

G

2

0

0

2

B

1

6

C

5

1

1

3

D

4

3

E

3

0

0

4

H

1

6

I

4

2

2

IDLE TIME = 3

EFFICIENCY = (1 - (SIGMA(IDLE TIME) / (NUMBER OF WORKSTATIONS * CYCLE TIME))) * 100

EFFICIENCY = (1 - (3 / (4 * 7))) * 100 = 89.29 OR 89%